# MedIAnomaly: A comparative study of anomaly detection in medical images

### v3

@@ Abstract
Anomaly detection (AD) aims at detecting abnormal samples that deviate from the expected normal patterns. Generally, it can be trained merely on normal data, without a requirement for abnormal samples, and thereby plays an important role in the recognition of rare diseases and health screening in the medical domain. Despite the emergence of numerous methods for medical AD, we observe a lack of a fair and comprehensive evaluation, which causes ambiguous conclusions and hinders the development of this field. To address this problem, this paper builds a benchmark with unified comparison. Seven medical datasets with five image modalities, including chest X-rays, brain MRIs, retinal fundus images, dermatoscopic images, and histopathology whole slide images, are curated for extensive evaluation. Thirty typical AD methods, including reconstruction and self-supervised learning-based methods, are involved in comparison of image-level anomaly classification and pixel-level anomaly segmentation. Furthermore, for the first time, we formally explore the effect of key components in existing methods, clearly revealing unresolved challenges and potential future directions. The datasets and code are available at https://github.com/caiyu6666/MedIAnomaly.
@@ 1 Introduction
Anomaly detection (AD) is a fundamental machine learning problem [25, 78], which aims to detect abnormal samples that deviate from expected normal patterns. It works under the assumption that a large number of normal samples with similar patterns are readily available, while abnormal samples, with diverse and unknown patterns, are difficult to collect comprehensively. Therefore, existing AD methods generally train the model merely on normal data. This alleviates the requirement for abnormal samples, facilitating the application of AD in the medical domain, especially in the recognition of rare diseases and health screening where abnormal samples are usually rare and diverse. Take the case of the United States, where more than 7,000 rare diseases affect more than 30 million people, though each disease alone affects fewer than 200,000 people (about 0.06% of the population). [^1] https://www.fda.gov/patients/rare-diseases-fda Their rarity and diversity make it infeasible to collect enough samples for each disease to train a typical supervised classification model. This fact makes AD an appropriate solution to assist radiologists in recognizing potential abnormalities from medical images.
@@
Numerous studies [94, 93, 68, 107, 95, 22] have been dedicated to medical AD. However, the lack of a comprehensive and fair comparison makes some conclusions unclear and hinders the development of this area. On the one hand, the utilization of different datasets or partitions in existing works impedes reproducibility and comparability, thereby undermining the reliability of the findings. Consequently, some studies underestimate the baseline and draw one-sided conclusions. For instance, several papers [109, 108, 124, 125] performed experiments on datasets such as Hyper-Kvasir [18] or OCT2017 [53]. However, as shown in Table 1, we observe that a simple auto-encoder (AE) implemented by us, with default settings described in Section 3.2.1, achieves nearly perfect performance on these datasets. This suggests that these datasets are not challenging for AD, and therefore evaluating AD methods on them is not suitable. On the other hand, despite plenty of methods belonging to the same paradigm, the absence of a unified implementation leads to different network architectures and diverse training tricks, resulting in unfair comparison. Take the case of reconstruction-based methods, where f-AnoGAN [93] adopts residual blocks [40] with multiple convolutional layers for the encoder and decoder, AE-Flow [125] chooses Wide ResNet-50-2 [120] as the encoder, while AE-U [68] uses plain convolutional layers without any specific designs. These different network architectures could interfere with the comparison of methodologies and make ambiguous results. Overall, the comparisons based on improper datasets or inconsistent implementations are inherently unreasonable. To address these issues, comprehensive datasets and a unified implementation of methodologies for medical AD are in high demand.
@@
    Through our experiments and analysis, we highlight unresolved challenges and potential future directions in the field of medical anomaly detection.
@@ 3.1.4 Summary of the datasets
Fig. 5 presents normal and abnormal examples from the seven datasets. Table 2 summarizes details of the image modality, supported task, and repartition in the datasets. Regaring the repartition, there is no validation set available for model selection due to the unavailability of abnormal samples and labels during the training of AD models.
@@ 3.2.1 Unified implementation of the reconstruction methods
To ensure fairness, we employ the same AE for all reconstruction-based methods whenever possible. The encoder consists of four convolutional blocks, each downsampling the resolution to $\frac{H_{in}}{2}\times\frac{W_{in}}{2}$. The encoder blocks are then followed by a flattening operation and two fully-connected layers, which compress the feature to a vector of dimension $d$. The decoder exhibits an architecture symmetric to the encoder. Table 3 provides a summary of our default configuration for this AE.
@@ 3.2.2 Unified implementation of the self-supervised methods
To ensure fairness, we employ ResNet18 [40], excluding classification layers, as the feature extractor for all SSL-based methods whenever possible. The input resolution is set to $224\times 224$. For additional specifics, we use the unified implementation from Schlüter et al. [95] as the framework for all one-stage segmentation approaches, and the implementation from https://github.com/Runinho/pytorch-cutpaste as the foundation for all two-stage approaches. Since most two-stage approaches utilize the classification task as the pretext task, we implement the one-stage classification approaches based on the same framework, but retain the classification head for evaluation. Table 4 provides a summary of our default configuration for SSL methods.
@@ 4 Experiments and analysis
To provide a comprehensive evaluation and analysis for medical AD methods, we assess a total of thirty AD methods on five different image modalities. This extensive coverage guarantees the generalizability of our results and findings. Table 5 provides a summary of the methods in our experiments. In addition, we evaluate the performance of directly using pixel intensity as the anomaly score [71], denoted as ‘Intensity’, which reflects the distinguishability between normal and abnormal pixel intensity within datasets.
Tables 6 and 7 present the performance of the methods on AnoCls and AnoSeg, respectively. Additionally, for qualitative analysis, we visualize reconstructed images and prediction maps generated by image-reconstruction methods in Fig. 6, and visualize prediction maps generated by other methods in Fig. 7. In the subsequent sections, we will first analyze the state-of-the-arts, and then delve into a detailed discussion of important findings from both the reconstruction and SSL methods.
@@ 4.1 Comparison of the state-of-the-arts
In the absence of the ImageNet weights, we notice that reconstruction methods outperform SSL methods in image-level AnoCls. When excluding the use of ImageNet weights, the SOTA SSL method is two-stage AnatPaste. However, this method exhibits unstable performance across datasets. It fails to outperform the vanilla AE on LAG and BraTS2021 datasets, indicating a high sensitivity to abnormal patterns. In contrast, AE-U and DAE, which are the SOTA reconstruction methods that do not use ImageNet weights, demonstrate stable and competitive performance on most datasets.
@@ 4.2.1 Effects of network architectures in AE
Overall, these experimental results demonstrate that current AE architectures for medical AD does not benefit from large resolutions or large networks. This perspective on reconstruction-based AD suggests that increasing the complexity of the model may be unnecessary. Consequently, it naturally raises a question of how the reconstruction methods effectively work for AD. This question will be further discussed in the following Section 4.2.2.
@@ 4.2.2 Effects of latent space restrictions in AE
To begin, we discuss latent dimension reduction. The results in Table 12 show that a remarkably compact latent size (e.g., 4–32) yields a substantial performance boost when applied to datasets containing local regional anomalies (i.e., RSNA, VinDr-CXR, Brain Tumor, LAG, and BraTS2021 datasets, as mentioned in Section 3.1.4). For example, decreasing the latent size from 128 to 4 improves the evaluation metrics by 12.4% AUC and 10.0% AP on RSNA dataset, 12.7% AUC and 12.2% AP on VinDr-CXR dataset, and 7.3% AUC and 6.9% AP on LAG dataset. In the case of brain MRI modalities, specifically the Brain Tumor and BraTS2021 datasets, a similar trend is observed, albeit with a larger optimal latent size of approximately 32 and 16, respectively.
@@
Furthermore, we explore the performance of the spatial latent size, where the bottleneck comprises $1\times 1$ convolutional layers instead of flattening and fully-connected layers. This allows the latent code to retain spatial shapes, such as $\mathbf{z}\in\mathbb{R}^{1\times 4\times 4}$ or $\mathbf{z}\in\mathbb{R}^{2\times 4\times 4}$. We observe that while the spatial AE significantly reduces the number of parameters compared to the dense AE, the performance may decrease on certain datasets. Specifically, both the dense AE with $\mathbf{z}\in\mathbb{R}^{16}$ and the spatial AE with $\mathbf{z}\in\mathbb{R}^{1\times 4\times 4}$ have the same number of latent variables. However, the dense AE ($\mathbf{z}\in\mathbb{R}^{16}$) has 2.35M parameters, whereas the spatial AE ($\mathbf{z}\in\mathbb{R}^{1\times 4\times 4}$) only has 0.22M parameters. Regarding performance, the spatial version suffers from a significant drop on Brain Tumor and LAG datasets. This phenomenon can be attributed to the fact that the spatial bottleneck compresses different regions to a consistent degree, regardless of the varying amounts of information present in these regions.
In addition to the simplest latent dimension reduction, there exist other popular designs that impose latent space restrictions in various ways, such as VAE [54], Constrained AE [29], and MemAE [36]. However, we observe from Table 6 that none of the existing approaches for restricting the latent space demonstrate superior performance compared to latent dimension reduction. Potential reasons are that VAE samples the latent code from the learnt distribution instead of using a deterministic vector, making the reconstruction indeterministic, while Constrained AE does not explicitly restrict the latent space. The unexpected failure of MemAE [36], despite its success on various natural images, raises questions. MemAE introduced a memory bank to store prototypes of normal latent features during training and utilized a combination of these prototypes to replace the latent feature of the input. It hypothesized a distinct difference between normal and abnormal features, aiming to render abnormal inputs unreconstructible using the combination of normal prototypes. However, this hypothesis is not necessarily valid due to the inherent nature of AE [13, 21]. Consequently, we attribute the failure of MemAE to the inseparability of latent features of normal and abnormal medical images. To substantiate our argument, we visualize the latent features of normal and abnormal medical images obtained from AE and present the performance of OC-SVM, which is built on these features, in Fig. 8. The results clearly demonstrate that these features are mixed together and inseparable by OC-SVM. This finding contradicts the underlying hypothesis of MemAE and provides support for our argument, ultimately explaining the failure of MemAE in medical scenarios.
@@ 4.2.5 Effects of noise types in DDPM
Current DDPM-based AD methods rely on the addition of noise during the diffusion process to corrupt abnormal regions, thereby facilitating pseudo-healthy reconstructions. Wyatt et al. [115] demonstrated the superiority of their proposed Simplex noise over the original Gaussian noise used in DDPM for corrupting abnormal regions, showcasing better performance on brain MRIs. In our experiments, we evaluate the performance of AnoDDPM [115] equipped with these two noise types on a wider range of datasets. Our findings align with their conclusions and provide additional insights into the limitations of DDPM-based AD methods stemming from their reliance on noise.
@@ 4.3.1 Methods without ImageNet weights
To explain the results, we conduct t-SNE visualization of representations learned through AnatPaste and CutPaste. To validate the effectiveness of realistic synthesis, we compare the visualization of the two typical synthesis approaches. As depicted in Fig. 10(a), a model trained through CutPaste can easily differentiate between the normal data and the unrealistic CutPaste samples. However, due to the significant discrepancy between real anomalies and synthetic ones, this model struggles to distinguish real anomalies from normal data. In contrast, Fig. 10(b) demonstrates that the model trained on realistic synthetic data can naturally recognize real anomalies by leveraging the classification training. This observation underscores the importance of the realism of synthetic data.
Additionally, to understand the superiority of the two-stage paradigm, let us consider the example provided in Fig. 10(a). The visualization highlights that although the classifier trained on unrealistic synthetic data fails to recognize real anomalies, the training process enables the model to effectively distinguish between the features of normal and abnormal samples. As a result, the two-stage paradigm, which removes the classification head and utilizes the learned representations, outperforms the one-stage paradigm, which directly employs the subpar classification head.
@@ 5.2 Latent space configuration of AE
Section 4.2.2 demonstrates that the latent space is the core component in reconstruction-based AD, with a simple adjustment of the latent dimensions proving to be the most effective means of controlling the latent space and achieving satisfactory performance. Unfortunately, existing approaches still rely on testing results to find the optimal configuration, leaving the question of how to adjust the latent space to its optimal state for diverse datasets unresolved. Recently, Cai et al. [21] leveraged information theory to reveal that the entropy of an optimal latent space should align with that of normal data. However, they did not develop a method to explicitly measure this entropy. To tackle this challenge, a promising direction would be to quantify the information entropy of normal training data, and develop self-adaptive methods that dynamically constrain the latent space to approach the entropy of normal data on different datasets. This approach would enable the determination of the theoretically optimal solution of AE-based AD methods.
@@ 6 Conclusion
This paper presents a comprehensive benchmark for medical anomaly detection, incorporating seven datasets and a comprehensive comparison of thirty typical methods. Our extensive evaluation and analysis reveal several key findings and challenges to guide future research in this field.

### v4

@@ Abstract
Anomaly detection (AD) aims at detecting abnormal samples that deviate from the expected normal patterns. Generally, it can be trained merely on normal data, without a requirement for abnormal samples, and thereby plays an important role in rare disease recognition and health screening in the medical domain. Despite the emergence of numerous methods for medical AD, the lack of a fair and comprehensive evaluation causes ambiguous conclusions and hinders the development of this field. To address this problem, this paper builds a benchmark with unified comparison. Seven medical datasets with five image modalities, including chest X-rays, brain MRIs, retinal fundus images, dermatoscopic images, and histopathology images, are curated for extensive evaluation. Thirty typical AD methods, including reconstruction and self-supervised learning-based methods, are involved in comparison of image-level anomaly classification and pixel-level anomaly segmentation. Furthermore, for the first time, we systematically investigate the effect of key components in existing methods, revealing unresolved challenges and potential future directions. The datasets and code are available at https://github.com/caiyu6666/MedIAnomaly.
@@ 1 Introduction
Anomaly detection (AD) is a fundamental machine learning problem [25, 78], which aims to detect abnormal samples that deviate from expected normal patterns. It works under the assumption that a large number of normal samples with similar patterns are readily available, while abnormal samples, with diverse and unknown patterns, are difficult to collect comprehensively. Therefore, existing AD methods generally train the model merely on normal data. This alleviates the requirement for abnormal samples, facilitating the application of AD in the medical domain, especially in rare disease recognition and health screening where abnormal samples are usually rare and diverse. Take the case of the United States, where over 7,000 rare diseases affect more than 30 million people, though each disease alone affects fewer than 200,000 people (about 0.06% of the population). [^1] https://www.fda.gov/patients/rare-diseases-fda Their rarity and diversity make it infeasible to collect enough samples for each disease to train a typical supervised model. This fact makes AD an appropriate solution to assist radiologists in recognizing potential abnormalities from medical images.
@@
Numerous studies [94, 93, 68, 107, 95, 22] have been dedicated to medical AD. However, the lack of a comprehensive and fair comparison makes some conclusions unclear and hinders the development of this area. On the one hand, the use of different datasets or partitions in existing studies impedes reproducibility and comparability. As a result, some studies underestimate the baseline and draw biased conclusions. For instance, several papers [109, 108, 124, 125] performed experiments on datasets such as Hyper-Kvasir [18] or OCT2017 [53]. However, as shown in Table 1, a simple auto-encoder (AE) implemented by us, with default settings described in Section 3.2.1, achieves nearly perfect performance on these datasets. This suggests that these datasets are not challenging for AD, and therefore evaluating AD methods on them is not suitable. On the other hand, despite many methods belonging to the same paradigm, the absence of a unified implementation leads to different network architectures and training tricks, resulting in unfair comparison. Take the case of reconstruction-based methods, where f-AnoGAN [93] adopts residual blocks [40] with multiple convolutional layers for the encoder and decoder, AE-Flow [125] chooses Wide ResNet-50-2 [120] as the encoder, while AE-U [68] uses plain convolutional layers without any specific designs. These differences could interfere with the comparison of methodologies and make ambiguous results. Overall, the comparisons based on improper datasets or inconsistent implementations are unreasonable. To address these issues, comprehensive datasets and a unified implementation of methodologies for medical AD are in high demand.
@@
    Through experiments and analysis, we highlight unresolved challenges and potential future directions in the field of medical anomaly detection.
@@ 3.1.4 Summary of the datasets
Fig. 5 presents normal and abnormal examples from the seven datasets. Table 2 summarizes details of the image modality, supported task, and repartition in the datasets. Regarding the repartition, there is no validation set available for model selection due to the unavailability of abnormal samples and labels during the training of AD models.
@@ 3.2.1 Unified implementation of the reconstruction methods
To ensure fairness, we employ the same AE for all reconstruction-based methods whenever possible. The encoder consists of four convolutional blocks, each downsampling the resolution to $\frac{H_{in}}{2}\times\frac{W_{in}}{2}$. The encoder blocks are then followed by a flattening operation and two fully-connected layers, which compress the feature to a vector of dimension $d$. The decoder exhibits an architecture symmetric to the encoder. Table 3 summarizes our default configuration for this AE.
@@ 3.2.2 Unified implementation of the self-supervised methods
To ensure fairness, we employ ResNet18 [40], excluding classification layers, as the feature extractor for all SSL-based methods whenever possible. The input resolution is set to $224\times 224$. For additional specifics, we use the unified implementation from Schlüter et al. [95] as the framework for all one-stage segmentation approaches, and the implementation from https://github.com/Runinho/pytorch-cutpaste as the foundation for all two-stage approaches. Since most two-stage approaches utilize the classification task as the pretext task, we implement the one-stage classification approaches based on the same framework, but retain the classification head for evaluation. Table 4 summarizes our default configuration for SSL methods.
@@ 4 Experiments and analysis
To provide a comprehensive evaluation and analysis for medical AD methods, we assess a total of thirty AD methods on five different image modalities. This extensive coverage guarantees the generalizability of our results and findings. In addition, we evaluate the performance of directly using pixel intensity as the anomaly score [71], denoted as ‘Intensity’, which reflects the distinguishability between normal and abnormal pixel intensity within datasets. Table 5 summarizes the methods in our experiments.
Tables 6 and 7 present the performance of the methods on AnoCls and AnoSeg, respectively. For qualitative analysis, we visualize reconstructed images and prediction maps generated by image-reconstruction methods in Fig. 6, and visualize prediction maps generated by other methods in Fig. 7. In the subsequent sections, we will first analyze the state-of-the-arts, and then delve into a detailed discussion of important findings from both the reconstruction and SSL methods.
@@ 4.1 Comparison of the state-of-the-arts
In the absence of the ImageNet weights, reconstruction methods outperform SSL methods in image-level AnoCls. When excluding the use of ImageNet weights, the SOTA SSL method is two-stage AnatPaste. However, this method exhibits unstable performance across datasets. It fails to outperform the vanilla AE on LAG and BraTS2021 datasets, indicating a high sensitivity to abnormal patterns. In contrast, AE-U and DAE, which are the SOTA reconstruction methods that do not use ImageNet weights, demonstrate stable and competitive performance on most datasets.
@@ 4.2.1 Effects of network architectures in AE
Overall, these experimental results demonstrate that current AE architectures for medical AD does not benefit from large resolutions or large networks. This perspective on reconstruction-based AD suggests that increasing the complexity of the model may be unnecessary. It naturally raises a question of how the reconstruction methods effectively work for AD. This question will be further discussed in the following Section 4.2.2.
@@ 4.2.2 Effects of latent space restrictions in AE
To begin, we discuss latent dimension reduction. Table 12 shows that a remarkably compact latent size (e.g., 4–32) yields a substantial performance boost when applied to datasets containing local regional anomalies (i.e., RSNA, VinDr-CXR, Brain Tumor, LAG, and BraTS2021 datasets, as mentioned in Section 3.1.4). For example, decreasing the latent size from 128 to 4 improves the evaluation metrics by 12.4% AUC and 10.0% AP on RSNA dataset, 12.7% AUC and 12.2% AP on VinDr-CXR dataset, and 7.3% AUC and 6.9% AP on LAG dataset. In the case of brain MRI modalities, specifically the Brain Tumor and BraTS2021 datasets, a similar trend is observed, albeit with a larger optimal latent size of approximately 32 and 16, respectively.
@@
Furthermore, we explore the performance of the spatial latent size, where the bottleneck comprises $1\times 1$ convolutional layers instead of flattening and fully-connected layers. This allows the latent code to retain spatial shapes, such as $\mathbf{z}\in\mathbb{R}^{1\times 4\times 4}$ or $\mathbf{z}\in\mathbb{R}^{2\times 4\times 4}$. While the spatial AE significantly reduces the number of parameters compared to the dense AE, the performance may decrease on certain datasets. Specifically, both the dense AE with $\mathbf{z}\in\mathbb{R}^{16}$ and the spatial AE with $\mathbf{z}\in\mathbb{R}^{1\times 4\times 4}$ have the same number of latent variables. However, the dense AE ($\mathbf{z}\in\mathbb{R}^{16}$) has 2.35M parameters, whereas the spatial AE ($\mathbf{z}\in\mathbb{R}^{1\times 4\times 4}$) only has 0.22M parameters. Regarding performance, the spatial version suffers from a significant drop on Brain Tumor and LAG datasets. This phenomenon can be attributed to the fact that the spatial bottleneck compresses different regions to a consistent degree, regardless of the varying amounts of information present in these regions.
In addition to the simplest latent dimension reduction, there exist other popular designs that impose latent space restrictions in various ways, such as VAE [54], Constrained AE [29], and MemAE [36]. However, Table 6 shows that none of the existing approaches for restricting the latent space demonstrate superior performance compared to latent dimension reduction. Potential reasons are that VAE samples the latent code from the learnt distribution instead of using a deterministic vector, making the reconstruction indeterministic, while Constrained AE does not explicitly restrict the latent space. The unexpected failure of MemAE [36], despite its success on various natural images, raises questions. MemAE introduced a memory bank to store prototypes of normal latent features during training and utilized a combination of these prototypes to replace the latent feature of the input. It hypothesized a distinct difference between normal and abnormal features, aiming to render abnormal inputs unreconstructible using the combination of normal prototypes. However, this hypothesis is not necessarily valid due to the inherent nature of AE [13, 21]. Consequently, we attribute the failure of MemAE to the inseparability of latent features of normal and abnormal medical images. To substantiate our argument, we visualize the latent features of normal and abnormal medical images obtained from AE and present the performance of OC-SVM built on these features, in Fig. 8. The results illustrate that these features are mixed together and inseparable by OC-SVM. This finding contradicts the underlying hypothesis of MemAE and provides support for our argument, ultimately explaining the failure of MemAE in medical scenarios.
@@ 4.2.5 Effects of noise types in DDPM
Current DDPM-based AD methods rely on the addition of noise during the diffusion process to corrupt abnormal regions, thereby facilitating pseudo-healthy reconstructions. Wyatt et al. [115] demonstrated the superiority of their proposed Simplex noise over the original Gaussian noise used in DDPM for corrupting abnormal regions, showcasing better performance on brain MRIs. We evaluate the performance of AnoDDPM [115] equipped with these two noise types on a wider range of datasets. Our findings align with their conclusions and provide additional insights into the limitations of DDPM-based AD methods stemming from their reliance on noise.
@@ 4.3.1 Methods without ImageNet weights
To elucidate the rationale behind these conclusions, we conduct t-SNE visualization of representations learned through AnatPaste and CutPaste. For the first conclusion (the efficacy of realistic synthesis), we compare the visualization of the two typical synthesis approaches. As depicted in Fig. 10(a), a model trained through CutPaste can easily differentiate between the normal data and the unrealistic CutPaste samples. However, due to the significant discrepancy between real anomalies and synthetic ones, this model struggles to distinguish real anomalies from normal data. In contrast, Fig. 10(b) illustrates that the model trained on realistic synthetic data can naturally recognize real anomalies by leveraging the classification training. This observation underscores the importance of the realism of synthetic data. For the second conclusion (the superiority of the two-stage paradigm), let us consider the example provided in Fig. 10(a). The visualization highlights that although the classifier trained on unrealistic synthetic data fails to recognize real anomalies, the training process enables the model to effectively distinguish between the features of normal and abnormal samples. As a result, the two-stage paradigm, which removes the classification head and utilizes the learned representations, outperforms the one-stage paradigm, which directly employs the subpar classification head.
@@ 5.2 Latent space configuration of AE
Section 4.2.2 demonstrates that the latent space is the core component in reconstruction-based AD, with a simple adjustment of the latent dimensions proving to be the most effective means of controlling the latent space and achieving satisfactory performance. Unfortunately, existing approaches still rely on testing results to find the optimal configuration, leaving the question of how to adjust the latent space to its optimal state for diverse datasets unresolved. Recently, Cai et al. [21] leveraged information theory to reveal that the entropy of an optimal latent space should align with that of normal data, but they did not develop a method to explicitly measure this entropy. To tackle this challenge, a promising direction would be to quantify the information entropy of normal training data and develop self-adaptive methods that dynamically constrain the latent space to approach the entropy of normal data on different datasets. This approach would enable the determination of the theoretically optimal solution of AE-based AD methods.
@@ 6 Conclusion
This paper presents a comprehensive benchmark for medical anomaly detection, incorporating seven datasets and a comparison of thirty typical methods. Our extensive evaluation and analysis reveal several key findings and challenges to guide future research in this field.